OpenStudy (cupcakehippo):

Help pls medal and fan :) solve for y x=(3-4y)^2

OpenStudy (pokieokie):

hmmmmm

OpenStudy (openstudier013087):

Square root both sides

OpenStudy (openstudier013087):

And then subtract 3

OpenStudy (cupcakehippo):

what do you mean by both sides

OpenStudy (openstudier013087):

It's because you have to get rid of the square so what you do to one side you do to the other

OpenStudy (openstudier013087):

To get rid of the square, square root it

OpenStudy (cupcakehippo):

then what will it be

OpenStudy (openstudier013087):

square root of x=3-4y

OpenStudy (pokieokie):

Thanks, im kinda in a rush before k12 closes out the school

OpenStudy (cupcakehippo):

okay then what

OpenStudy (openstudier013087):

Then subtract 3 from both sides so it will be square root of x -3=-4y

OpenStudy (cupcakehippo):

why subtract 3

OpenStudy (cupcakehippo):

woudnt you add 4 or someting

OpenStudy (openstudier013087):

No you can't because you have to divide by 4

OpenStudy (openstudier013087):

4 is attached to the y so to separate them you have to divide

OpenStudy (cupcakehippo):

so what do i subtract 3 by...the x... and 3?

OpenStudy (openstudier013087):

Yes and remember it is square root of x

OpenStudy (cupcakehippo):

then what

OpenStudy (openstudier013087):

\[\sqrt{x}-3=-4y\]

OpenStudy (openstudier013087):

Then you divide each side by -4

OpenStudy (cupcakehippo):

then what will i get

OpenStudy (cupcakehippo):

dont we have to get rid of the x

OpenStudy (openstudier013087):

Is this multiple choice?

OpenStudy (cupcakehippo):

no

OpenStudy (openstudier013087):

This can go both ways either they want you to solve for y and just leave the x alone or they want you to solve for the y like we just did and replace x with (3-4y)^2

OpenStudy (cupcakehippo):

...

OpenStudy (openstudier013087):

In this section what are you learning?

OpenStudy (openstudier013087):

Are you solving for actual values for the variables?